OpenStudy (anonymous):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. my equation is f(x)=2(x-6)^2+4

OpenStudy (anonymous):

\[f(x)=2(x-6)^2+4\] right?

OpenStudy (anonymous):

\[f(7)=2(7-6)^2+4=6\]

OpenStudy (anonymous):

\[f(4)=2(4-6)^2+4=12\] if my arithmetic is correct

OpenStudy (anonymous):

your final job is to compute \[\frac{f(7)-f(4)}{7-3}=\frac{6-12}{4}=...\]

OpenStudy (anonymous):

so it would be \[\frac{ 6 }{ 4 }\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sorry im really bad at math, hahaha

OpenStudy (anonymous):

since \(6-12=-6\) it would be \(-\frac{6}{4}=-\frac{3}{2}\)

OpenStudy (anonymous):

watch them signs!

OpenStudy (anonymous):

OH, because a smaller number is subtracting to a larger so it would end up being negative! i get it lol